A 5kg block was placed on an inclined plane(50 degrees with respect to the vertical).Forces A is at angle C with respect to the horizontal. Force B is parallel with the incline.The coefficients of static and kenetic friction are 0.3 and 0.2 respectively.
A=197KN
B=49 KN
C=7 degrees
DETERMINE IF THERE IS MOTION AND ITS DIRECTION
1
Expert's answer
2020-12-29T15:32:50-0500

"N=mg\cos40°+197000\cdot\sin47°=5\cdot9.8\cdot\cos40°+197000\cdot\sin47°="


"=144114(N)"


"49000+mg\sin40°-197000\cos47°=49000+5\cdot9.8\cdot\sin40°-197000\cos47°="


"=-85322(N)"


"F_f=0.3N=0.3\cdot144114=43234(N)"


So, "F_{total}=-85322+43234=-42088(N)"


The block will move up the inclined plane. Answer